OpenStudy (anonymous):

16^3 = 4^x Please show work

OpenStudy (anonymous):

false?

OpenStudy (anonymous):

this can also be written as 16^3 = 16^1, asking if the two sides are equal. well it's false because 16^3 > 16^1

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

sorry got confuse, the equation goes like this: 16^3 = 4^x

OpenStudy (anonymous):

16 = 4^2 4^(3*2) = 4^x 4^6 = 4^x

OpenStudy (theviper):

x=6

OpenStudy (anonymous):

Thanks for everything!

OpenStudy (theviper):

yw:)

OpenStudy (theviper):

LOL but I have written only two words